OpenStudy (anonymous):

helpp Given the equation 3x - 4y = 8, which equation below would cause a consistent-dependent system? 3x + 4y = -8 6x - 8y = 12 9x - 12y = 24 16x + 12y = -10

OpenStudy (anonymous):

am so confused can you walk me through it a bit

OpenStudy (anonymous):

yea sort of,like i understand the example you gave above but am not sure if my answer to my question is correct i belive the answer is C. is that correct?

OpenStudy (anonymous):

oh thank you so much,it was actrally your example that helped me cause i just switched the numbers

OpenStudy (anonymous):

can you see if you can help me with a different question?

OpenStudy (anonymous):

How many solutions will this system of equations have? x - 2y = 24 3x - 6y = 72 *one solution *two solution *no solution *infinite solution

OpenStudy (anonymous):

I know that its not a two solution or infinate solution so am stuck between 1 or no solution

OpenStudy (anonymous):

oh but how did you know it was the same line

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

oh womyou explain it better then my teacher lol

OpenStudy (anonymous):

*wow

OpenStudy (anonymous):

can you hep me with just one more

OpenStudy (anonymous):

ok Determine which system below will produce infinitely many solutions. *2x + 5y = 24 2x + 5y = 42 * 3x - 2y = 15 6x + 5y = 11 *4x - 3y = 9 -8x + 6y = -18 * 5x - 3y = 16 -2x + 3y = -7 i think its C but am not sure